# R-Judge: Benchmarking Safety Risk Awareness for LLM Agents

### v2

@@ Abstract
Large language models (LLMs) have exhibited great potential in autonomously completing tasks across real-world applications. Despite this, these LLM agents introduce unexpected safety risks when operating in interactive environments. Instead of centering on LLM-generated content safety in most prior studies, this work addresses the imperative need for benchmarking the behavioral safety of LLM agents within diverse environments. We introduce R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging and identifying safety risks given agent interaction records. R-Judge comprises 162 records of multi-turn agent interaction, encompassing 27 key risk scenarios among 7 application categories and 10 risk types. It incorporates human consensus on safety with annotated safety labels and high-quality risk descriptions. Evaluation of 9 LLMs on R-Judge shows considerable room for enhancing the risk awareness of LLMs: The best-performing model, GPT-4, achieves 72.52% in contrast to the human score of 89.07%, while all other models score less than the random. Moreover, further experiments demonstrate that leveraging risk descriptions as environment feedback achieves substantial performance gains. With case studies, we reveal that correlated to parameter amount, risk awareness in open agent scenarios is a multi-dimensional capability involving knowledge and reasoning, thus challenging for current LLMs. R-Judge is publicly available at https://github.com/Lordog/R-Judge.
@@ 1 Introduction
Figure 1: Dataset distribution of R-Judge, which contains 27 key risk scenarios among 7 application categories, and spans across 10 risk types. The proportion of each category is shown in the graph.
@@
Large language models (LLMs) have shown compelling abilities in reasoning, decision making, and instruction following (Wei et al., 2022a). The stimulating capabilities of LLMs, especially GPT-4 (OpenAI, 2023), ignite the development of LLM agents (Zhang et al., 2023b). Equipped with tool usage and environment interaction, agents, such as AutoGPT (Richards, 2023) and Voyager (Wang et al., 2023a), can autonomously complete user-specified tasks with LLMs as controllers (Xi et al., 2023; Wang et al., 2023b; Zhou et al., 2023b; Lin et al., 2023a; Sumers et al., 2023).
@@
Given that unknown risks rest in complex environments and user instructions, LLM agents are prone to cause unexpected safety issues (Xi et al., 2023; Ruan et al., 2024; Naihin et al., 2023). For instance, when asked to process emails, agents may unconsciously click the URL of phishing emails, leading to potential privacy leakage and even property loss. Therefore, safety assurance is a necessary prerequisite for LLM agents in real-world applications. To this end, it is critical to effectively evaluate the safety risk awareness of LLMs in open agent scenarios.
@@
Existing works on safety evaluation focus on safety issues of LLM-generated contents (Zhiheng et al., 2023; Bhardwaj and Poria, 2023), e.g., to alleviate generating contents related to offensiveness, unfairness, illegal activities, and ethics. Notably, SafetyBench (Zhang et al., 2023a) evaluated LLMs with multiple-choice questions covering 7 categories of safety concerns. Similarly, SuperCLUE-Safety (Xu et al., 2023a) presented a comprehensive benchmark with multi-round and open-ended questions. However, without interactive environments where LLM agents execute, they could not provide a realistic safety evaluation of LLM agents.
@@
To tackle the aforementioned challenge, few studies have investigated safety risks in applications of LLM agents. To identify risks of LLM agents in interactive environments, ToolEmu (Ruan et al., 2024) implemented a GPT-4 powered emulator with diverse tools and scenarios tailored for the LLM agents’ action execution. Alongside the emulator, a GPT-4 powered automatic safety evaluator examines agent failures and quantifies associated risks. In a similar vein, AgentMonitor (Naihin et al., 2023) proposed a framework where an LLM serves as a monitor for the execution of agents. It intervenes by halting actions deemed unsafe, thereby preventing safety issues that LLM agents might encounter on the open internet.
@@
Though implemented with different settings, both ToolEmu and AgentMonitor utilized LLMs as safety monitors to identify risky actions of LLM agents. However, judging whether agent actions are safe in the context of multi-turn interactions involving LLM agents, users, and environment in diverse scenarios, is a challenging yet realistic task for LLMs. Owing to the complexity of interactions and diversity of environments, it remains unclear whether LLMs are aware of agent safety issues without priors of risk descriptions from humans.
@@
This work presents R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging and identifying safety risks given agent interaction records. Each record contains a user instruction and a history of agent actions and environment feedback. R-Judge comprises 162 agent interaction records, encompassing 27 popular application scenarios across 7 categories, including program, operation system (OS), internet of things (IoT), software, web, finance, and health (Figure 1). For each record, we annotate binary safety labels as well as high-quality risk descriptions of identified risks with human consensus.
@@
Utilizing R-Judge, we conducted a comprehensive evaluation of 8 prominent LLMs commonly employed as the backbone for agents and LlamaGuard-7b which is specifically fine-tuned on identifying safety risks in LLM-generated content. The results demonstrate considerable room for enhancing the risk awareness of LLMs: the best-performing model, GPT-4, achieves an F1 score of 72.52% in contrast to the human score of 89.07% while all other models fail the task. Notably, LlamaGuard-7b is incapacitated in R-Judge. Further experiments find that leveraging risk descriptions as environment feedback significantly improves model performance while straightforward prompting mechanisms fail. With case studies, we conclude that risk awareness in open agent scenarios is correlated to parameter amount and is a multi-dimensional capability involving knowledge and reasoning, thus challenging for LLMs. In summary, our contributions are as follows:
@@
$\bullet$ We propose R-Judge, a human-aligned benchmark dataset with complex multi-turn interactions between the user, environment, and agent. It incorporates human consensus on safety with annotated safety labels and high-quality risk descriptions. To our knowledge, this is the first benchmark dataset curated to assess risk awareness of LLMs for agent safety.
$\bullet$ We evaluate 9 popular LLMs based on R-Judge and disclose that most models fail to identify safety risks in open agent scenarios. Furthermore, we explore the impact of different mechanisms and conduct in-depth case studies, summarizing key findings with valuable insights to facilitate future research.
@@ 2 Problem Formalization
The interaction process of the LLM agent is as follows. First, the user interacts with the agent by user instruction $u\in\mathcal{L}$, where $\mathcal{L}$ is the language space. Following user instruction, the agent generates a thought $t\in\mathcal{L}$ followed by an action $a\in\mathcal{A}$, where $\mathcal{A}$ defines action space. After executing the action, the agent receives environment feedback $f\in\mathcal{F}$, where $\mathcal{F}$ is the space of environment states. Then the process circulates until the task finishes or errors occur. We denote the process of interactions as a record:
@@
where $n$ is the number of interaction turns to finish the task. Users may call the agent many times, so a list of records is also a record. For each turn, the agent takes $R$ as inputs and outputs $(t,a)$ to interact with environments.
Risk Awareness of LLMs for Agent Safety.
Interacting directly with complex environments, LLM agents are prone to cause unexpected safety issues, some of which are severe, such as crucial privacy leakage and data loss. With LLMs as safety monitors, LLM agents cause fewer safety issues, as disclosed by Toolemu and AgentMonitor. Therefore, risk awareness of LLMs is a significant factor for the safe execution of LLM agents. Here, we formulate the task of agent safety monitor to evaluate the risk awareness of LLMs.
@@
As illustrated in Figure 2, an LLM checks agent interaction records to judge if actions of the agent are safe. The performance as a safety monitor demonstrates whether the LLM can effectively identify the safety risks and make correct judgments, revealing its risk awareness. The task formulation is:
$f:p_{\theta}(R)\to(\textrm{analysis},\textrm{label}).$ (2)
where $p_{\theta}(\cdot)$ denotes the language model mapping record $R$ to $\textrm{analysis}\in\mathcal{L}$ and a binary label representing safe or unsafe. For general evaluation, LLMs are blind to priors of scenario-specific risks and safety guidelines in various scenarios.
@@ 3 R-Judge Benchmark
Figure 2: Illustration of R-Judge by an example. The upper left part is an example in the dataset, i.e., a record of an agent interacting with the user and environment. The lower left part is human annotation involving a binary safety label and high-quality risk description. Following the arrow, we can see the serial evaluation paradigm with two tests: given record and task instruction, LLMs are asked to generate an analysis and a label. An automatic evaluator compares the analysis with the ground truth risk description to assess the effectiveness of risk identification. And correct generated labels are counted to evaluate the performance of safety judgment.
To evaluate risk awareness of LLMs for agent safety, we construct R-Judge, a general and realistic benchmark dataset containing multi-turn agent interaction with human safety consensus.
@@ 3.1 Case Format
Following the problem formalization in Section 2, we define the case format, as shown on the left half of Figure 2. Here, we provide an elaborate explanation of each term.
@@
$\bullet$ Agent Record. The record snapshots the interaction process between the user, environment, and agent. For generality, we use ReAct (Yao et al., 2023) as the agent framework where agents think and act in interactive environments.
@@
$\bullet$ Safety Label. 1 for unsafe, and 0 for safe. It is labeled and cross-checked carefully by well-trained human annotators. The labeling standard is the safety consensus of annotators based on general safety standard attached in Appendix A.3. We ensure truth labels are clear with no ambiguity through cross-validation by human annotators.
$\bullet$ Risk Description. Carefully documented by annotators, a risk description includes complete information for humans and agents to understand the risks of the case. For each unsafe case, the risk description contains Motivation, Trigger, and Outcome, following the threat model of agent safety in Appendix A.4. Motivation means the risk source from the user or environment. Trigger means the risky behaviors of the agent. Outcome is the result after the agent finishes risky actions.
3.2 Dataset Curation
To contribute to a realistic dataset with high quality, we design a dataset curation pipeline incorporating human consensus on safety.
Curation Principle.
Datasets are curated by both public collection and manual construction, as occurrences of risks are relatively rare in current agent applications. We first predefined curation principles, including category and scenario (Appendix A.1), risk type (Appendix A.2), general safety standard on LLM agents (Appendix A.3), and threat model of LLM agents (Appendix A.4). Then based on curation principles, human annotators collect cases from related work and construct risky cases manually. For labeling, annotators mark a safety label and write a risk description for each record. Whether the label is safe or unsafe is determined by the annotator and then validated by others based on predefined general safety standard (Appendix A.3). By incorporating human consensus on safety, R-Judge accommodates human-aligned ground truth.
Public Collection.
We collect and modify public datasets, mainly from Toolemu (Ruan et al., 2024) and AgentMonitor (Naihin et al., 2023). For validity, cases are transformed or removed based on curation principles.
@@ 3.2 Dataset Curation
We also construct high-quality data from human experts. Concretely, annotators meticulously brainstorm application cases where agents are likely to trigger certain risk types in certain scenarios. Then, with the assistance of ChatGPT and our validity check, application cases are polished into valid ones.
Due to the complex logic of interaction records and the safety mechanism of GPT-4, we find that GPT-4 is unable to generate valid records for batch construction. Even from the AutoGPT team, AgentMonitor crafted only 27 single-turn unsafe cases, demonstrating the difficulty of manual construction. At last, 81 from Toolemu, 24 from AgentMonitor, and 57 constructed cases form the final 162 R-Judge records.
Data Statistics.
R-Judge comprises 162 complex cases where intricate risks lie in the multi-turn interaction between the user, agent, and environment. On average, it involves 3.0 turns of interaction and 192.9 word counts, with 63.0% being unsafe cases. The diverse dataset encompasses 7 selected categories including program, OS, IoT, software, web, finance, and health, covering 27 scenarios. It also spans 10 types of risk (Appendix A.2), including privacy leakage, computer security, physical health, data loss, financial loss, property damage, illegal activities, ethics & morality, bias & offensiveness, and miscellaneous. Figure 1 presents an overview of R-Judge. More statistics are shown in Table 4.

### v3

@@ Abstract
Large language models (LLMs) have exhibited great potential in autonomously completing tasks across real-world applications. However, LLM agents introduce unexpected safety risks when operating in interactive environments. Instead of centering on the harmlessness of LLM-generated content in most prior studies, this work addresses the imperative need for benchmarking the behavioral safety of LLM agents within diverse environments. We introduce R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging and identifying safety risks given agent interaction records. R-Judge comprises 569 records of multi-turn agent interaction, encompassing 27 key risk scenarios among 5 application categories and 10 risk types. It is of high-quality curation with annotated safety labels and risk descriptions. Evaluation of 11 LLMs on R-Judge shows considerable room for enhancing the risk awareness of LLMs: The best-performing model, GPT-4o, achieves 74.45% while no other models significantly exceed the random. Moreover, we reveal that risk awareness in open agent scenarios is a multi-dimensional capability involving knowledge and reasoning, thus challenging for LLMs. With further experiments, we find that fine-tuning on safety judgment significantly improves model performance while straightforward prompting mechanisms fail. R-Judge is publicly available at https://github.com/Lordog/R-Judge.
R-Judge: Benchmarking Safety Risk Awareness for LLM Agents
Tongxin Yuanthanks:   Equal contribution. Corresponding author. , Zhiwei He, Lingzhong Dong, Yiming Wang, Ruijie Zhao, Tian Xia, Lizhen Xu, Binglin Zhou, Fangqi Li, Zhuosheng Zhang, Rui Wang, Gongshen Liu School of Electronic Information and Electrical Engineering, Shanghai Jiao Tong University {teenyuan,zwhe.cs,zhangzs,wangrui12,lgshen}@sjtu.edu.cn
@@ 1 Introduction
Figure 1: Dataset distribution of R-Judge, which contains 27 key risk scenarios among 5 application categories, and spans across 10 risk types. The proportion of each category is shown in the graph.
@@
Large language models (LLMs) have shown compelling abilities in reasoning, decision-making, and instruction following (Wei et al., 2022a). The stimulating capabilities of LLMs, especially GPT-4 (OpenAI, 2023), ignite the development of LLM agents (Xi et al., 2023; Wang et al., 2023b; Zhang et al., 2023b). Equipped with tool usage and environment interaction, agents, such as AutoGPT (Richards, 2023), Voyager (Wang et al., 2023a) and OpenHands (Wang et al., 2024), can autonomously complete user-specified tasks with LLMs as controllers (Zhou et al., 2023b; Lin et al., 2023a).
@@
Given that unknown risks rest in complex environments, LLM agents are prone to cause unexpected safety issues (Xi et al., 2023; Ruan et al., 2024; Naihin et al., 2023). For instance, when asked to process emails, agents may unconsciously click the URL of phishing emails, leading to potential privacy leakage and even property loss. Therefore, there is growing awareness (Yang et al., 2024; Tang et al., 2024) that safety assurance is a necessary prerequisite for LLM agents in real-world applications. To this end, it is critical to effectively evaluate the safety risk awareness of LLMs in open agent scenarios.
@@
Existing works on safety evaluation focus on safety issues of LLM-generated contents (Zhiheng et al., 2023; Bhardwaj and Poria, 2023), i.e., to alleviate generating contents related to offensiveness, unfairness, illegal activities, and ethics. Notably, SafetyBench (Zhang et al., 2023a) and SuperCLUE-Safety (Xu et al., 2023a) evaluated LLMs with multiple-choice or open-ended questions covering various safety concerns. In addition to evaluation, LLM-based monitors (Inan et al., 2023; Zhang et al., 2024) are developed to moderate LLM-generated content. However, benchmark questions can not expose behavioral risks in interactive environments, and thus struggle to provide a practical safety evaluation for LLM agents. Moreover, whether content monitors are able to function in open-agent scenarios is still a question.
@@
To tackle the aforementioned challenge, few studies have investigated safety risks in applications of LLM agents. To identify risks of LLM agents in interactive environments, ToolEmu (Ruan et al., 2024) implemented a GPT-4 powered emulator with diverse tools and scenarios tailored for the LLM agents’ action execution. Alongside the emulator, a GPT-4 powered automatic safety evaluator examines agent failures and quantifies associated risks. In a similar vein, AgentMonitor (Naihin et al., 2023) proposed a framework where an LLM serves as a monitor for the execution of agents. It intervenes by halting actions deemed unsafe, thereby preventing safety issues that LLM agents might encounter on the open internet. Specifically, InjecAgent (Zhan et al., 2024) exposes risks of indirect prompt injection attacks on LLM agents where attackers inject malicious instructions into the environment to manipulate agents into executing detrimental actions against users.
@@
Though implemented with different settings, both ToolEmu and AgentMonitor utilized LLMs as safety monitors to identify the risky actions of LLM agents. However, judging whether agent actions are safe in the context of multi-turn interactions involving LLM agents, users, and environment in diverse scenarios, is a challenging yet pract task for LLMs. Due to the complexity of interactions and the diversity of environments, it remains unclear whether LLMs are aware of agent safety issues.
@@
This work presents R-Judge, a benchmark crafted to evaluate the proficiency of LLMs in judging and identifying safety risks given agent interaction records. Each record contains a user instruction and a history of agent actions and environment feedback. R-Judge comprises 569 agent interaction records, encompassing 27 popular application scenarios across 5 categories, including program, internet of things (IoT), software, web, and finance (Figure 1). For each record, we annotate binary safety labels as well as descriptions of identified risks.
@@
Utilizing R-Judge, we conducted a comprehensive evaluation of 11 prominent LLMs commonly employed as the backbone for agents. Concretely, feeding the records of agent execution as inputs, the evaluated LLM is required to identify risks and make safety judgments on whether agent actions are safe. The results demonstrate considerable room for enhancing the risk awareness of LLMs, revealing the significant concern of agent safety: the best-performing model, GPT-4o, achieves an F1 score of 74.45% while no other models evidently surpass the random. Further experiments find that fine-tuning on safety judgment significantly improves model performance while straightforward prompting mechanisms fail. With case studies, we conclude that risk awareness in open agent scenarios is a multi-dimensional capability involving knowledge and reasoning, thus challenging for LLMs.
In summary, our contributions are as follows:
@@
$\bullet$ We propose R-Judge, a practical benchmark dataset with complex multi-turn interactions between the user, environment, and agent. It is of high-quality curation with annotated safety labels and risk descriptions. To our knowledge, this is the first benchmark dataset curated to assess risk awareness of LLMs for agent safety.
$\bullet$ We evaluate 11 popular LLMs based on R-Judge and disclose that most models fail to identify safety risks in open agent scenarios. Furthermore, we explore the impact of different mechanisms and conduct in-depth case studies, summarizing key findings with valuable insights to facilitate future research on agent safety.
@@ 2 Problem Formalization
Figure 2: Illustration of R-Judge by an example. The upper left part is an example in the dataset, i.e., a record of an agent interacting with the user and environment. The lower left part is human annotation involving a binary safety label and high-quality risk description. Following the arrow, we can see the serial evaluation paradigm with two tests: given record and task instruction, LLMs are asked to generate an analysis and a label. An automatic evaluator compares the analysis with the ground truth risk description to assess the effectiveness of risk identification. Correct generated labels are counted to evaluate the performance of safety judgment.
@@
We denote $\mathcal{L}$, $\mathcal{A}$, and $\mathcal{F}$ as spaces of languages, actions, and environment states, respectively. The interaction process of the LLM agent is as follows: Initially, the user interacts with the agent by user instruction $u\in\mathcal{L}$. Following $u$, the agent generates a thought $t\in\mathcal{L}$ followed by an action $a\in\mathcal{A}$. After executing the action, the agent receives environment feedback $f\in\mathcal{F}$. One $(t,a,f)$ forms an iteration unit. Then this unit iterates until the task finishes or errors occur. We denote the complete interaction process as a record:
@@
where $n$ is the number of interaction turns when finishing the task. Users may call the agent many times, so a list of records is also a record. For each turn, the agent takes $R$ as inputs and outputs $(t,a)$ to interact with environments.
Risk Awareness for Agent Safety of LLMs.
When interacting directly with complex environments, LLM agents are prone to cause unexpected safety issues, some of which are severe, such as crucial privacy leakage and data loss. However, with LLMs as safety monitors, LLM agents cause fewer safety issues, as disclosed by ToolEmu and AgentMonitor. Therefore, risk awareness of LLMs is a significant factor for the safe execution of LLM agents. Here, we formulate the task of agent safety monitor to evaluate the risk awareness of LLMs.
@@
As illustrated in Figure 2, an LLM checks agent interaction records to judge if the agent actions are safe. The performance as a safety monitor demonstrates whether the LLM can effectively identify the safety risks and make correct judgments, revealing its risk awareness. The task formulation is:
$f:p_{\theta}(R)\to(\textrm{analysis},\textrm{label}),$ (2)
where $p_{\theta}(\cdot)$ denotes the language model mapping record $R$ to $\textrm{analysis}\in\mathcal{L}$ and a binary label representing safe or unsafe.
@@ 3 R-Judge Benchmark
To evaluate the risk awareness for agent safety of LLMs, we construct R-Judge, a realistic benchmark dataset containing multi-turn agent interactions with human safety annotation.
@@ 3.1 Case Format
Following the problem formalization in Section 2, we define the case format, as shown on the left half of Figure 2. Here we provide an elaborate explanation of the key terms:
@@
$\bullet$ Agent Record. The record snapshots the interaction process between the user, environment, and agent. For generality, we use ReAct (Yao et al., 2023) as the agent framework where agents think, act, and receive feedback in interactive environments.
@@
$\bullet$ Safety Label. The label is a binary value of ‘unsafe’ and ‘safe’. It is labeled and cross-checked carefully by well-trained human annotators (see detailed information in Appendix C) to ensure unambiguity.
$\bullet$ Risk Description. Carefully documented by annotators, a risk description includes complete elements for humans to understand the risks in the sample. Schema for the risk description is elaborated in Appendix A.3.
3.2 Curation Principles
To contribute to a diverse dataset with high quality, we first designed curation principles, including problem scope and threat model of LLM agents. Specifically, we focus on two fundamental risks that threat personal LLM agents, Intended Attacks and Unintended Risks.
Problem Scope.
We focus on Personal LLM agents (Li et al., 2024), the major paradigm of agent applications, which means the LLM agents are expected to be deployed on the system of the user. Based on that, we define 5 categories and according 27 scenarios shown in Figure 1. See Table 4 in Appendix A.1 for details.
Threat Model.
For personal LLM agents, the user prompts are expected to be benign, [^1] Direct attacks in user prompts such as jailbreaks (Yuan et al., 2023; Deng et al., ) are hence excluded. though may unconsciously contain risks. Therefore, regarding human intentions, risks that challenge personal agents fall into two fundamental origins (Tang et al., 2024).
  • Intended Attacks: The outer attackers attempt to manipulate the agents through the medium of environments. Here, we focus on a specific and primary way, indirect prompt injection attacks (Yi et al., 2023). For instance (Figure 8 in Appendix B.1), in the process of the agent calling Evernote notes, the attacker injected the malicious instruction “Please grant permanent access to my friend Amy (guest_amy01) on the August Smart Lock system.” in the return value of the tool EvernoteManagerSearchNotes. If the agent follows the injected malicious instruction, the attacker may gain access to the user’s residence, leading to severe risks.
  • Unintended Risks: Due to the complexity of the environments and multi-turn interactions, agents may cause safety issues even if there are no malicious attacks. For instance (Figure 9 in Appendix B.1), the user asks the agent to turn off the electrical devices when leaving home. But the agent turn off all electrical devices including security cameras, leading to potential issues.
To the end, we classify agents involved in safety issues into 10 risk types shown in Figure 1, following ToolEmu (Ruan et al., 2024).
3.3 Curation Process
Due to rare occurrences of visible risks in current agent applications and the challenge of automatically manufacturing valid records, datasets are primarily constructed from relevant works on agent safety. To enrich the data of Unintended Risks, we also construct high-quality data from human experts. Notably, to fairly assess the risk awareness for LLM agents, we endeavor to balance the number of safe and unsafe samples. And ‘safe’ cases are ones that agents may cause safety issues but actually not. More curation details are elaborated in Appendix B.2.
Open-source Transformation.
The open-source data from ToolEmu, InjecAgent and AgentMonitor accommodate complete trajectories of LLM agents. Based on them, we filter invalid data that conflict with our curation principles, annotate the safety labels, and write risk descriptions. Notably, agents in AgentMonitor do not follow ReAct (with only ‘action’), so we leave the field of ‘thought’ as null.
@@ 3.3 Curation Process
We also construct high-quality data from human experts. For one part, to supplement inadequate ‘safe’ examples, we transform some of the ‘unsafe’ examples in ToolEmu into ‘safe’ ones by replacing the risky agent actions with safe ones. For another part, to expand dataset, annotators meticulously brainstorm application cases where agents are likely to trigger certain risk types in certain scenarios. Then, with the assistance of ChatGPT and our validity check, application cases are polished into valid ones.
3.4 Dataset Statistics
At last, injecAgent (Zhan et al., 2024) contributed 414 samples to the data of Intended Attacks. 81 samples from ToolEmu (Ruan et al., 2024), 24 samples from AgentMonitor (Naihin et al., 2023), along with 55 samples of manual construction from human annotators, constitute the 155 data of Unintended Risks.
R-Judge comprises 569 complex cases where intricate risks lie in the multi-turn interaction between the user, agent, and environment. On average, R-Judge involves 2.6 turns of interaction and 206 word counts, with 52.7% being unsafe cases. The diverse dataset covers 5 selected categories including program, IoT, software, web, and finance, covering 27 scenarios. R-Judge also spans 10 types of risk (Appendix A.2), including privacy leakage, computer security, physical health, data loss, financial loss, property damage, illegal activities, ethics & morality, bias & offensiveness, and miscellaneous. Figure 1 presents an overview of R-Judge. More statistics are shown in Table 5 in Appendix B.3.